OpenStudy (skylyerbailey):

Halp :) Which ordered pairs in the form (x, y) are solutions to the equation 4x – 5y = 24? Choose all answers that are correct. A. (−9, −12) B. (1, −4) C. (4, 8) D. (6, 0)

OpenStudy (kamibug):

HAI SKY ! <3 GURL DIDYA TRY PLUGGIN IN THE OPTIONS FOR X AND Y ??? ;DDD

OpenStudy (skylyerbailey):

HAI KAMI <333333 I think the answer is all except C

OpenStudy (anonymous):

yep

OpenStudy (skylyerbailey):

Thanks @987....Kami is that right?

OpenStudy (kamibug):

Let's try em out ! 4(-9) - 5(-12) = 24 -36 - (-60) = 24 - 36 + 60 = 24 A. IS ONE YEAH GURL<3 4(1) - 5(-4) = 24 4 + 20 = 24 YAY B. IS ANOTHER 4(4) - 5(8) = 24 16 - 40 = 24 UH OH C ISNT RIGHT ITD BE NEGATIVE 24 :PPP 4(6) - 5(0) = 24 24 - 0 = 24 AND YEAHHH D. IS ALSO CORRECT YAY ! YOU DID IT GURL<333

OpenStudy (anonymous):

yay

OpenStudy (skylyerbailey):

HAHAHA @KamiBug you are so funny X'DDD thanks!!